OpenStudy (anonymous):

an individual saves $5000 a bank account at the beginning of each year for 10 years. No further saving or withdrawals are made from the account. determines the total amount saved if the annual interest rate is 8% compounded. (a) annually (b) semi-annually

OpenStudy (anonymous):

i think its annualy

OpenStudy (anonymous):

i should calculate both of them

OpenStudy (anonymous):

5000(1+8/100)^10 ?

OpenStudy (tkhunny):

Really? This s NOT different than those e did before IF you understad teh basic principles. i = 0.08 RA = 1 + i = 1.08 -- Annual Accumulation Factor RS = 1 + i/2 = 1.04 -- Semi-Annual Accumulation Factor Annual Accumulation \(5000(RA^{10} + RA^{9} + ... + RA)\) -- Done. Semi-Annual Accumulation \(5000(RS^{20} + RS^{18} + ... + RS^{2})\) -- Done. Seriously, forget the formulas and learn to draw the map!

OpenStudy (anonymous):

what grade is this?

OpenStudy (anonymous):

annual it is 5000(1.08^10-1/1.08-1) =7243

OpenStudy (anonymous):

first year at university

OpenStudy (anonymous):

oops then i shouldnt be helping :)

OpenStudy (anonymous):

annual sorry it is 72432 and semi annual 5000(1.04^20-1/1.04-1)=148890

OpenStudy (tkhunny):

No good. You are thinking about formulas and not about basic principles. If you think about formulas, then every chapter or seciont will seem new and will be increasingly confusing. If you think about basic principles, you will never need t obe confused again. \(5000(RA^{10} + RA^{9} + ... + RA) = 5000\frac{RA - RA^{11}}{1-RA}\) There are other wayso write this formula, b none more useful if yojust created it yourself and you have ultimate confiden that it correct.

OpenStudy (anonymous):

ohh kk 10 + first term kk understand!

OpenStudy (tkhunny):

You have to start thinking about it more. You seem to be substantially guessing. For starters, accumulating 5,000 annually for 10 years is AT LEAST 50,000. Do not EVER accept a values less than 50,000. Your "7243" cannot possibly be even close. It should make sense before you think it is the answer.

OpenStudy (anonymous):

annual =7190 but you can calculate the semi annual day ? if it start from (r^2 >r^20

OpenStudy (anonymous):

how it can 5000 time 10 if you have interest 8%?

OpenStudy (tkhunny):

72432 is no good. Are you rounding badly again? 7232 is no good. You are making payments at the END of the year. It is NOT 5000*10 = 50000. This is just a reality check. If you get 7243, that clearly is wrong. The answer MUST be greater than 50,000.

OpenStudy (tkhunny):

You're still just guessing. One step at a time. i = 0.08 -- Annual interest rate r = 1+i = 1.08 -- Annual accumulation factor "5000 at the beginning of each of 10 years" #1 - 5000*r^10 #2 - 5000*r^9 #3 - 5000*r^8 ... #10 - 5000*r^1 Add them up: 5000*r^10 + 5000*r^9 + 5000*r^8 + ... + 5000*r^1 A little algebra: 5000*(r^10 + r^9 + r^8 + ... + r^1) A little algebra: \(5000\dfrac{r - r^{11}}{1-r}\) A little calculator work r^11 = 2.3316389971 r - r^11 = -1.2516389971 1 - r = -.08 Final answer: 5000((-1.2516389971)/(-0.08)) = 5000*15.6454874638 = 78227.44 There is no need to guess.

OpenStudy (tkhunny):

Calculator: Type = Beginning of Period N = 10 i = 0.08 or 8.00 -- However your calculator expects to see this value. Pmt = 5000 or maybe -5000 -- Again, make sure you understand what your calculator wants. PV = 0 Calculate FV

OpenStudy (anonymous):

i just confusing because i did R^11-1/0.8*5000 i did for annual because i thought that first term should 1.08^0 =1> 1.08^10

OpenStudy (tkhunny):

We need to back up a little bit. Have you studied the "Order of Operations"?

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

It is usually very early on - possibly before algebra. Are you sure you haven't seen it? It usually comes out like "PEMDAS".

OpenStudy (anonymous):

no usually interest coumpound and geomectric serie only

OpenStudy (tkhunny):

That's okay. It is not as prominent in all cultures. U.S. math texts usually harp on it. It is a way to write more clearly in a standardized fashion so that the things you write can be more frequently and more clearly understood. For example, when you write "R^11-1/0.8*5000 ", how do I know the order in which you did this calculation? Usually, some judicious parentheses will clear it up. (R^11-1)/0.8*5000 Sometimes spacing makes it more clear. (R^11 - 1)/0.8*5000 Sometimes order makes it more clear 5000*(R^11 - 1)/0.8 See what I mean?

OpenStudy (anonymous):

yes know i understand we in Lebanon that you should see how we studied now i understand that

OpenStudy (anonymous):

1.33163/0.8*5000=83227

OpenStudy (tkhunny):

Good. This is just encouragement to try to write more clearly. Make sure it says what you want it to say. Remember that we are not inside your head. -) Okay, enough of that. Back to the problem... Your statement "5000*(R^11 - 1)/0.8" has a couple of problems. 1) You have 0.8 in the denominator. No good. Maybe 0.08? 2) You cannot have R^11 with 1 in the numerator and the interest rate in the denominator unless you have 11 payments. You have only 10. 3) All this leads me to believe you are still kind of just guessing. You're not quite sure which formula to use. How are we doing?

OpenStudy (tkhunny):

"1.33163/0.8*5000=83227" This s a good example of what I was talking about. You WROTE "0.8" but you MEANT and USED "0.08". Write what you mean!

OpenStudy (anonymous):

look what i did r+R^1.....r^10 so 11 term i did by my head calculation 8/100=0.8 i see it that it is wrong it is 0.08 so that why i did 5000(r^11-1=r^0)/0.08 that i a did

OpenStudy (tkhunny):

No. r + R^1 <== What does that mean? You have an extra term.

OpenStudy (anonymous):

ahh kk so it r^10-1?/0.08

OpenStudy (anonymous):

if it 10 year r^10-1/0.08*5000

OpenStudy (tkhunny):

Closer, but still no. That would be correct for payments at the END of the period. You have payments at the BEGINNING of each period.

OpenStudy (anonymous):

hard to understand term begenning and end period ^^

OpenStudy (tkhunny):

BEGINNING just has one more accumulation period than END. End: \(\dfrac{r^{N} - 1}{i}\) Beginning: \(r\cdot\dfrac{r^{N} - 1}{i} = \dfrac{r^{N+1} - r}{i} = \dfrac{r^{N} - 1}{i\cdot v} = \dfrac{r^{N} - 1}{d}\) You may see any or all of these other forms.

OpenStudy (anonymous):

d? what is it

OpenStudy (anonymous):

i*v ?

OpenStudy (tkhunny):

It is a standard definition. \(v =\dfrac{1}{1+i}\)

OpenStudy (anonymous):

we didnt take standard definition

OpenStudy (tkhunny):

Another standard definition \(d = i\cdot v = \dfrac{i}{1+i}\) You should be familiar with these. You WILL encounter them.

OpenStudy (anonymous):

but we didnt studied yet i can use standar for my homework...

OpenStudy (tkhunny):

You can use what you can define.

OpenStudy (tkhunny):

The semi-annual problem is almost the same. You just have to redefine r and i. For annual, we had r = 1+i = 1.08 For semi-annual, still with annual payments, we have \(r = \left(1 + \dfrac{0.08}{2}\right)^2\). This makes i = 0.0816 for the semianual calulation.

OpenStudy (anonymous):

kk that make send sens i try to understand i how find annual anwser without using standar formual we didnt learned

OpenStudy (anonymous):

annual i found 5000(1.08)(1.08^10-1/1.08-1)

OpenStudy (tkhunny):

I believe that, except for the lack of clarity. I would have preferred 5000(1.08)(1.08^10 - 1)/(1.08-1) You should be able to solve the semi-annual one using i = 0.0816

OpenStudy (anonymous):

5000(1.04)(1.04^20-1)/(1.04-1)

OpenStudy (tkhunny):

No. That would be good for 10 years of semi-annual payments. That is NOT what you have. You have 10 years of ANNUAL payments that are compounded semi-annually.

OpenStudy (anonymous):

semie annual so 5000(1+8/2/100)^10*2

OpenStudy (anonymous):

when we use geometrci serie it should like i did !

OpenStudy (tkhunny):

You need the equivalent annual interest related to your semi-annual compounding. Once you have that, it is EXACTLY like the first problem in this pair.

OpenStudy (tkhunny):

One ALWAYS uses the Geometric Series, otherwise the theory doesn't exist.. Do not separate what you are seeing from the Geometric Series fundamental development.

OpenStudy (anonymous):

mmmm...

OpenStudy (tkhunny):

Your compounding period is different from your payment period. It takes a little more thought. Remember this? \(i = 0.08\) \(RS = 1 + i/2 = 1.04 \) \(5000(RS^{20} +RS^{18} +...+ RS^{2})\) -- Done. \(5000\cdot\dfrac{RS^{20} - 1}{RS^{2} - 1}\cdot RS^{2}\) -- Also done.

OpenStudy (anonymous):

it should be rs^2? because it semi anual but why not 5000(1.04)?

OpenStudy (tkhunny):

Payments are only every other compounding period. There is not a payment every ½ year.

OpenStudy (anonymous):

ohhh kk thank you so much

OpenStudy (tkhunny):

Time 0 5000 Time ½ 5000*(1.04) Time 1 (5000*1.04*1.04 + 5000)

OpenStudy (anonymous):

understsand^^ finaly